# Revealing the impact of synthetic native samples and multi-tasking strategies in Hindi-English code-mixed humour and sarcasm detection

### v1

@@ Abstract
In this paper, we reported our experiments with various strategies to improve code-mixed humour and sarcasm detection. We did all of our experiments for Hindi-English code-mixed scenario, as we have the linguistic expertise for the same. We experimented with three approaches, namely (i) native sample mixing, (ii) multi-task learning (MTL), and (iii) prompting very large multilingual language models (VMLMs). In native sample mixing, we added monolingual task samples in code-mixed training sets. In MTL learning, we relied on native and code-mixed samples of a semantically related task (hate detection in our case). Finally, in our third approach, we evaluated the efficacy of VMLMs via few-shot context prompting. Some interesting findings we got are (i) adding native samples improved humor (raising the F1-score up to 6.76%) and sarcasm (raising the F1-score up to 8.64%) detection, (ii) training MLMs in an MTL framework boosted performance for both humour (raising the F1-score up to 10.67%) and sarcasm (increment up to 12.35% in F1-score) detection, and (iii) prompting VMLMs couldn’t outperform the other approaches. Finally, our ablation studies and error analysis discovered the cases where our model is yet to improve. We provided our code for reproducibility.
@@
Debajyoti Mazumder, Aakash Kumarthanks: This work was done during his stay at Indian Institute of Science Education and Research, Bhopal., Jasabanta Patro Intelligent Systems’ Lab (ISL), Department of Data Science and Engineering Indian Institute of Science Education and Research, Bhopal, India {debajyoti22, aakashk19, jpatro}@iiserb.ac.in
@@ 1 Introduction:
Humour and sarcasm are complex and subjective emotions that impact the nature of human communication. They can appear in different forms such as exaggeration, dark humour, gross humour, adult or slang expression, insult, offense, etc. Frenda et al. (2018); Ahuja (2019). Past study Bleakley and Sailofsky (2023) highlighted how they can affect politics amid tragedy. Thus, detecting these emotions is essential. It is also challenging, especially in a code-mixed setting, where two languages are mixed in a single sentence. This is because models now need to understand the context expressed through language alteration and detect humour and sarcasm in it. Code-mixing is a linguistic phenomenon common across multilingual speakers. Multilingual speakers are believed to outnumber monolingual speakers globally Tucker (1999). A large chunck of population in Asia, Europe, and North America know more then one language, this allows them to communicate among themselves by switching languages in a single utterance. It is more prominent in informal communications like social media posts and voice mails Patro et al. (2017). An example of humorous and sarcastic expression in Hindi-English code-mixed language is given in the following. More examples are presented in Figure 1 of Appendix B. In the following example, the English parts are marked in red and Hindi parts are marked in blue. We have provided translations for redability.
@@
The NLP community has shown significant interest in humour and sarcasm detection, mostly in monolingual settings Abulaish et al. (2020); Joshi et al. (2017, 2020). Unfortunately, there is relatively less focus on code-mixed humour and sarcasm Singh and Sharma (2023); Elayan et al. (2022); Chen et al. (2024a). Further, very few publicly annotated corpora are available in code-mixed settings Sitaram et al. (2019). The same is true for many low-resource languages as well. It acts as a significant bottleneck for research on this topic. The evolution of multilingual large language models (MLMs hereafter) opened an opportunity to improve the situation via cross-lingual learning. MLMs can now learn task-specific knowledge from a training corpus in one language and predict it for samples in a different language. MLMs can do this as they learn their embeddings from corpora spanning multiple languages. Cross-lingual learning has disadvantages as well. They can perform better only when training and test corpora share similar linguistic and cultural contexts Bigoulaeva et al. (2021); Gupta et al. (2022). Mazumder et al. (2024) hypothesized that since code-mixed corpora (in Hindi-English code-mixed language) and the native monolingual corpora (in Hindi and English languages) are likely to share similar linguistic and cultural context, the addition of native task samples in code-mixed training sets can improve code-mixed task performance. A set of empirical experiments has shown that adding Hindi and English hate samples in code-mixed hate training corpora improves code-mixed hate detection Mazumder et al. (2024). However, this hypothesis has yet to be tested in code-mixed Humour and Sarcasm detection. A detailed discussion of prior works in this direction is presented in the Appendix A. Further, we observed that Hindi humour and sarcasm datasets are not publicly available in Hindi-English code-mixed settings. Thus, blindly adding only English samples to code-mixed training sets may not improve the performance. In the present work, we proposed a multi-tasking strategy that works in such cases. We conducted various experiments based on three broad research questions. All of our experiments were done for Hindi-English code-mixed scenarios as we have the necessary linguistic expertise for them. It is essential to do the ablation studies and error analysis (presented in the Appendix). The three research questions are,
@@
R3: Do adding native samples in the prompting context of Very large MLMs (VMLMs hereafter) improve the performance?
@@
  • We integrated related tasks like hate detection with native samples in a multitask learning framework (refer Exp. 2: Section 3.3). Instead of sequential training, the model processed batches containing mixed-task samples. We conducted an ablation study to understand the role of the gating mechanism in the MTL framework (in Appendix F.1). This approach gave significant performance improvements for MLMs, with F1-score up to 10.67% increment in humour and 12.35% increment in sarcasm ($p<0.05$). With a gating mechanism, they better handled shorter contexts and misspelt samples.
  • We compared the performance of VMLMs using native and code-mixed examples as few-shots (refer Exp. 3: Section 3.4). However, neither type of example could improve the VMLM predictions.
@@ 2 Datasets:
English HASOC8 2261 3591 76.47% 60.68% 1.182 72% (overlap)
Table 1: Dataset statistics. Notation: # for number of samples, H denotes proportion of samples containing hurtful (offensive, aggressive and hateful) keywords from positive or negative class, IAA for Inter-Annotator Agreement and KL for symmetrized smoothed Kullback-Leibler divergence between word distributions of positive ($+ve$) and negative ($-ve$) samples.
@@ 3 Experiments:
In this section, we reported the details of our experiments conducted as a part of this study. Apart from reproducing the baselines (section 3.1), we designed three experiments each based on a unique research philosophy. The three philosophies are i) native sample mixing, ii) multi-task learning and iii) prompting very large multilingual language models (VMLMs). While the native sample mixing strategy intends to improve the code-mixed tasks by adding monolingual native samples to the code-mixed training sets, the multi-task learning strategy tries to do the same by learning linguistic knowledge from the samples of a third task. Our last strategy, i.e. prompting VMLMs, evaluates the performance of very large multilingual language models for the considered code-mixed tasks in a few-shot context prompting scenario. The detailed experimental set up is given in Appendix D.1.
@@ 3.1 Baselines:
Our first experiment explored the impact of native language samples after adding them to the code-mixed training sets. Past study Mazumder et al. (2024) reported that this strategy works for code-mixed hate detection. However, no one tested it for code-mixed humour and sarcasm detection. Further, in our case, even though there are publicly available English humour and sarcasm datasets (Table 1), we couldn’t find any Hindi datasets for the same. Thus, we choose to create silver annotated datasets by translating some portion of English datasets to Hindi using Google Translator API [^1] https://cloud.google.com/translate?hl=en. From the methodological point of view, we considered two types of models.
  • Statistical classifiers: We considered three statistical classifiers i.e. Naive Bayes (NB), Random Forest (RF), and Support Vector Machine (SVM) in our study. NB is known to perform better when there is high KL divergence between classes. We utilized word-level unigrams, bigrams, and trigrams as features. Past works showed these features to perform best Khandelwal et al. (2018); Swami et al. (1805).
@@
In our second approach, we explored the efficiency of multi-task framework to detect code-mixed humour and sarcasm by learning linguistic knowledge from the samples of a third task, i.e., here it is, hate detection. We chose hate detection as a third task because (i) it is semantically related to humour and sarcasm, (ii) samples of humour and sarcasm datasets contained hateful keywords (refer Table 1), and (iii) we could found Hindi, English and code-mixed samples available for this task. Our multi-task framework is inspired by the framework proposed by Rotman and Reichart (2022). They utilized a BERT-like architecture divided into two parts. The bottom module consists of eight lower layers of a transformer-based language model like BERT were common to the participating tasks. The top module, containing top four layers, were separately present for the individual tasks. Apart from that there is an additional top module present for parameter sharing. A gating mechanism connects all of the task-specific top modules with the additional top module. Authors experimented this framework on tasks like dependency parsing and named entity recognition and found that it performs better. In our case (i) we added a regularization term for soft parameter sharing of the final layers from the top module, and (ii) we froze the parameters of the bottom module during fine-tuning. The architecture of our framework is shown in Figure 2. We used embeddings from three widely used MLMs, i.e., (i) mBERT, (ii) XLM-R and (iii) MuRIL to initialize the layers of our MTL framework. We conducted an ablation study (refer Appendix F.1) to examine the role of the gating mechanism within the MTL models by removing the gate and comparing the performance with models that included the gating mechanism.
@@
In our third experiment, we evaluated the performance of very large multilingual language models in detecting humour and sarcasm in code-mixed texts. As their name suggests, they have a lot of parameters, thus, they are designed to work with prompting approaches rather than fine-tuning them. We utilized four VMLMs: (i) Gemma Team et al. (2024), (ii) Aya Expanse Üstün et al. (2024), (iii) Llama 3.1 Dubey et al. (2024) and (iv) GPT-4 Achiam et al. (2023) [^2] https://chatgpt.com/. These VMLMs were chosen for their strong performance in both Indic languagesWatts et al. (2024) and English. We used few-shot prompting scenario. Here, we show a few training examples in the prompt while asking the VMLMs to classify for a test sample. This strategy has proven its superiority in sentiment-related code-mixed tasksYadav et al. (2024). The prompt template and some of the examples are reported in Appendix E. We conducted many variants of this experiment by considering samples from native and code-mixed training set as few shots examples in the prompt.
@@ 4.1 Baseline results:
Table 2: Baselines results. Notation: originally reported scores are mentioned with ‘##\##’ mark and our reproduced results are marked as ‘†’.
@@
XLM-R $\underline{0.81}^{\#}$ ($\underline{0.83}^{\ast}$) 0.79 ($\underline{0.83}^{\ast}$) $0.81^{\#}$ $0.81^{\#}$ $0.81^{\#}$ 0.80 (0.83) $0.81^{\#}$
@@
In this section, we reported our observations from the first experiment. The F1-scores obtained from all models over the considered datasets and training scenarios are reported in Table 3. The F1-scores of best-performing models for the individual training scenarios (column-wise in Table 3) were marked in bold, while the second-best results are underlined. Similarly, the best-performing scenarios giving the highest $F1$ scores for individual models (row-wise in Table 3) were kept inside the parenthesis, while the second best scores are marked with ‘##\##’ superscript. The highest score for both tasks across all training scenarios and models are marked in blue. Following are our takeaways,
  • When trained with only code-mixed samples, mBERT gave the highest F1-score of 0.78 for humour detection, followed by XLM-R and MuRIL with an F1-score of 0.75 each. Similarly, for sarcasm detection, MuRIL reported the highest F1-score of 0.83 followed by IndicBERT and XLM-R, with an F1-score of 0.81 each.
  • On adding native samples to the code-mixed training sets, the Naive Bayes classifier showed no significant improvement in code-mixed Humour detection. The F1-scores still remained in the range of 0.74 - 0.75. On the other hand, For sarcasm detection, the performances declined sharply to the range of 0.32 - 0.42, compared to 0.74 when models were trained with only code-mixed samples. RF models have shown a performance decline for both code-mixed humour (decrease upto 0.04 in F1-score) and sarcasm (decrease upto 0.26 in F1-score) detection, except when synthetic Hindi samples from the HaHa dataset were added to the code-mixed training set. The improvement is (an increase of 0.01 in F1-score), however, not statistically significant. SVM models, too, have registered a decline in performance for both code-mixed humour (up to a decline of 0.07 in F1 scores) and sarcasm (up to a decline of 0.15 in F1 scores) detection after native samples were mixed with the code-mixed training sets.
  • Among the MLMs, after native sample mixing, we saw a performance decline for mBERT (up to 0.13 decrease in F1-score) in code-mixed humour detection. On the other hand, it showed significant improvement with an F1-score of 0.84 in detecting sarcasm, compared to 0.80 when trained with code-mixed samples. Similarly, for XLM-R we observed that the F1 score remained in the range of 0.72 - 0.76 and 0.79 - 0.83 for code-mixed humor and sarcasm detection after native sample mixing, compared to 0.75 and 0.81 respectively when trained with code-mixed samples alone. The slight improvements are however not statistically significant. In case of MuRIL, we saw many instances where it resulted statistically significant (p < 0.05) improvement (up to 0.04 and 0.06 raise in F1-score for humour and sarcasm detection respectively) after native sample mixing. Finally, we observed that IndicBERT model performed best after native sample mixing with an F1 scores of 0.79 and 0.88 in code-mixed humour and sarcasm detection, compared to 0.74 and 0.81 respectively (both statistically significant with p< 0.05) when trained with code-mixed samples.
@@
In this section, we presented our observations from the second experiment. The F1-scores obtained from all models for the considered datasets and training scenarios are reported in Table LABEL:tab:MTL_humor_results. The best-performing F1-scores for individual models (column-wise in Table LABEL:tab:MTL_humor_results) across all training scenarios were marked in bold, and the second best were underlined. Similarly, the F1-scores of best performing models for the individual training scenarios (row-wise in Table LABEL:tab:MTL_humor_results) were kept in parenthesis, while the second best were marked with ‘##\##’ superscript. The overall best scores were marked in red. Following were our takeaways,
[…]
  • For humor detection, $\text{mBERT}_{MTL}$ improved slightly and the F1 scores remained in the range 0.67 - 0.80, compared to 0.78 when trained with only code-mixed samples. However, the improvements were not statistically significant. On the other hand, for sarcasm detection, $\text{mBERT}_{MTL}$ resulted with a highest F1 score of 0.86, compared to 0.80 when trained with only code-mixed samples. The improvement is statistically significant.
[…]
  • Finally, $\text{MuRIL}_{MTL}$ reported the highest F1 score for humor detection, i.e., 0.83. The improvement is statistically significant, compared to 0.75 when trained on code-mixed samples. Similarly, for sarcasm detection, F1 scores ranged from 0.81 to 0.88 using $\text{MuRIL}_{MTL}$. The improved score of 0.88 is statistically significant ($p<0.05$) with respect to 0.83, when trained on code-mixed samples only.
  • Upon analyzing the improvements deeply, we found that the MTL models performed better for samples with shorter context length (refer to Figure 5). We reported the insights from the ablation study in Appendix F.1.
@@
In this section, we reported our observations from the third experiment. Out of the range of number of few shots listed in Table 6, the VMLMs achieved the highest F1-scores with eight few-shot examples. The F1-scores obtained from all models are reported in Table 4. The F1-scores of best-performing models for the individual prompting scenarios (column-wise in Table 4) were marked in bold, while the second-best results are underlined. Similarly, the best-performing F1-scores for individual models (row-wise in Table 4) were kept inside the parenthesis, while the second best scores are marked with ‘##\##’ superscript. The overall highest scores for both of the tasks are marked in blue. Following were our takeaways,
@@
Gemma (0.09) 0.04 $0.07^{\#}$ 0.02 0.04
Aya Expanse (0.73) (0.73) $\underline{0.72}^{\#}$ 0.67 (0.73)
Llama-3.1 (0.75) 0.68 $\textbf{0.73}^{\#}$ 0.55 0.64
GPT-4 (0.74) 0.56 $0.62^{\#}$ 0.56 0.47

### v2

@@ Abstract
In this paper, we reported our experiments with various strategies to improve code-mixed humour and sarcasm detection. Particularly, we tried three approaches: (i) native sample mixing, (ii) multi-task learning (MTL), and (iii) prompting and instruction finetuning very large multilingual language models (VMLMs). In native sample mixing, we added monolingual task samples to code-mixed training sets. In MTL learning, we relied on native and code-mixed samples of a semantically related task (hate detection in our case). Finally, in our third approach, we evaluated the efficacy of VMLMs via few-shot context prompting and instruction finetuning. Some interesting findings we got are (i) adding native samples improved humor (raising the F1-score up to 6.76%) and sarcasm (raising the F1-score up to 8.64%) detection, (ii) training MLMs in an MTL framework boosted performance for both humour (raising the F1-score up to 10.67%) and sarcasm (increment up to 12.35% in F1-score) detection, and (iii) prompting and instruction finetuning VMLMs couldn’t outperform the other approaches. Finally, our ablation studies and error analysis discovered the cases where our model is yet to improve. We provided our code [^1] https://github.com/islnlp/code-mix-humor-sarcasm-detection-EMNLP-2025 for reproducibility.
@@
Debajyoti Mazumder, Aakash Kumarthanks: This work was done during his stay at Indian Institute of Science Education and Research, Bhopal., Jasabanta Patro Department of Data Science and Engineering Indian Institute of Science Education and Research, Bhopal, India {debajyoti22, aakashk19, jpatro}@iiserb.ac.in
@@ 1 Introduction:
Humour and sarcasm are complex and subjective emotions that impact the nature of human communication. They can appear in different forms such as exaggeration, dark humour, gross humour, adult or slang expression, insult, offence, etc. Frenda et al. (2018); Ahuja (2019). Past study Bleakley and Sailofsky (2023) highlighted how they can affect politics amid tragedy. Detecting Humour and Sarcasm becomes more challenging in a code-mixed setting. This is because models now need to understand humour and sarcasm in an utterance expressed through altering multiple languages. More details on the phenomenon of code-mixing are presented in the Appendix. An example of humorous and sarcastic expression in Hindi-English code-mixed language is given in the following. More examples are presented in Figure 1 of Appendix B. In the following example, the English parts are marked in red, and the Hindi parts are marked in blue. We have provided their translations for readability.
@@
The NLP community has shown significant interest in monolingual humour and sarcasm detection Abulaish et al. (2020); Joshi et al. (2017, 2020). Unfortunately, there is relatively less focus on the code-mixed settings Singh and Sharma (2023); Elayan et al. (2022); Chen et al. (2024a). Therefore, we have few publicly annotated code-mixed corpora available, further acting as a bottleneck in developing new models Sitaram et al. (2019); Doğruöz et al. (2021); Winata et al. (2023). The evolution of multilingual large language models (MLM hereafter) has shown a new path to address this issue. They can learn task-specific knowledge from samples in one language and make predictions for samples in different languages. This phenomenon is known as cross-lingual learning. It is very effective if training and testing samples share similar linguistic and cultural contexts Bigoulaeva et al. (2021); Gupta et al. (2022). MLMs learn their embeddings from a corpora spanning multiple languages, thus they are aware of vocabulary of multiple languages. In the context of code-mixed languages, it means we can fine-tune the MLMs using native monolingual task samples (e.g. for Hindi and English task samples for a Hindi-English code-mixed task) and do prediction for code-mixed samples. The hypothesis is that since code-mixed corpora (in Hindi-English code-mixed language) and the native monolingual corpora (in Hindi and English languages) are likely to share similar linguistic and cultural contexts, the training with native task samples or adding them in code-mixed training sets can improve code-mixed task performance. In fact, Mazumder et al. (2024), in a set of empirical experiments, have shown that adding Hindi and English hate samples in code-mixed hate training corpora improves code-mixed hate detection Mazumder et al. (2024). However, nobody has tested it for code-mixed Humour and Sarcasm detection. A detailed discussion of prior works in this direction is presented in the Appendix A. Further, we observed that Hindi humour and sarcasm datasets are not publicly available. We experimented with synthetic Hindi samples and multi-tasking strategies to fill this gap. Overall, we asked for three research questions,
@@
R3: Do adding native samples in the prompting context or in instruction finetuning of Very large MLMs (VMLMs hereafter) improve the performance?
@@
  • We integrated a related task which is hate detection with native samples in a multitask learning framework (refer Exp. 2: Section 3.3). Instead of sequential training, the model processed batches containing mixed-task samples. We conducted an ablation study to understand the role of the gating mechanism in the MTL framework (in Appendix F). This approach gave significant performance improvements for MLMs, with F1-score up to 10.67% increment in humour and 12.35% increment in sarcasm ($p<0.05$). With a gating mechanism, they better handled shorter contexts and misspelt samples.
  • We compared the performance of VMLMs using native and code-mixed examples as few-shots (refer Exp. 3: Section 3.4). In the second set-up, we compared the VMLMs’ performance when native samples are addded to the code-mixed training set. However, neither type of set-ups could improve the VMLM predictions.
@@ 2 Datasets:
English HASOC9 2261 3591 76.47% 60.68% 1.182 72% (overlap)
Table 1: Dataset statistics. Notation: # for number of samples, H denotes proportion of samples containing hurtful (offensive, aggressive and hateful) keywords from positive or negative class, IAA for Inter-Annotator Agreement and KL for symmetrized smoothed Kullback-Leibler divergence between word distributions of positive (+ve+ve) and negative (ve-ve) samples.
@@ 3 Experiments:
In this section, we reported the details of our experiments conducted as a part of this study. Apart from reproducing the baselines (section 3.1), we designed three experiments each based on a unique research philosophy. The three philosophies are i) native sample mixing (section 3.2), ii) multi-task learning (section 3.3) and iii) prompting and instruction finetuning very large multilingual language models (VMLMs) (section 3.4). While the native sample mixing strategy intends to improve the code-mixed tasks by adding monolingual native samples to the code-mixed training sets, the multi-task learning strategy tries to do the same by learning linguistic knowledge from the samples of a third task (here, it is hate detection). Our last strategy, i.e. prompting and instruction finetuning VMLMs, evaluates the performance of very large multilingual language models for the considered code-mixed tasks in a few-shot context prompting and instruction finetuning scenarios. The detailed experimental set up is given in Appendix E.1.
@@ 3.1 Baselines:
Our first experiment explored the impact of native language samples after adding them to the code-mixed training sets. Past study Mazumder et al. (2024) reported that this strategy works for code-mixed hate detection. However, no one tested it for code-mixed humour and sarcasm detection. Further, in our case, even though there are publicly available English humour and sarcasm datasets (Table 1), we couldn’t find any Hindi datasets for the same. Thus, we choose to create silver annotated datasets by translating some portions of English datasets into Hindi using Google Translator API [^2] https://cloud.google.com/translate?hl=en. Note that accurately translating humour and sarcasm samples is still an open research topic. Thus, we used the most popular publicly available translation tool of current time. From the methodological point of view, we considered two types of models.
  • Statistical classifiers: We considered three statistical classifiers, i.e. Naive Bayes (NB), Random Forest (RF), and Support Vector Machine (SVM) in our study. NB is known to perform better when there is high KL divergence between classes. We utilized word-level unigrams, bigrams, and trigrams as features. Past works showed these features to perform best Khandelwal et al. (2018); Swami et al. (2018).
@@
In our second approach, we explored the efficiency of multi-task framework to detect code-mixed humour and sarcasm by learning linguistic knowledge from the samples of a third task, i.e., here it is, hate detection. We chose hate detection as a third task because (i) it is semantically related to humour and sarcasm, (ii) samples of humour and sarcasm datasets contained hateful keywords (refer Table 1), and (iii) we could found Hindi, English and code-mixed samples available for this task. Our multi-task framework is inspired by the framework proposed by Rotman and Reichart (2022). They utilized a BERT-like architecture divided into two parts. The bottom module consists of eight lower layers of a transformer-based language model like BERT were common to the participating tasks. The top module, containing top four layers, were separately present for the individual tasks. Apart from that there is an additional top module present for parameter sharing. A gating mechanism connects all of the task-specific top modules with the additional top module. Authors experimented this framework on tasks like dependency parsing and named entity recognition and found that it performs better. In our case (i) we added a regularization term for soft parameter sharing of the final layers from the top module, and (ii) we froze the parameters of the bottom module during fine-tuning. The architecture of our framework is shown in Figure 3. We used embeddings from three widely used MLMs, i.e., (i) mBERT, (ii) XLM-R and (iii) MuRIL to initialize the layers of our MTL framework. We conducted an ablation study (refer Appendix F) to examine the role of the gating mechanism within the MTL models by removing the gate and comparing the performance with models that included the gating mechanism.
@@
In our third experiment, we evaluated the performance of very large multilingual language models in detecting humour and sarcasm in code-mixed texts. As their name suggests, they have a lot of parameters, thus, they are designed to work well with prompting approaches rather than fine-tuning them fully. We utilized four VMLMs: (i) Gemma Team et al. (2024), (ii) Aya Expanse Üstün et al. (2024), (iii) Llama 3.1 Dubey et al. (2024) and (iv) GPT-4333https://chatgpt.com/ Achiam et al. (2023). These VMLMs were chosen for their strong performance in both Indic languages Watts et al. (2024) and English. We used two scenarios: (i) few-shot prompting and (ii) instruction finetuning using LoRA Hu et al. (2021) adapter. Here in the few shot set-up, we show a few training examples in the prompt while asking the VMLMs to classify for a test sample. This strategy has proven its superiority in sentiment-related code-mixed tasksYadav et al. (2024). The prompt template and some of the examples are reported in Appendix D.1. We conducted many variants of this experiment by considering samples from native and code-mixed training set as few shots examples in the prompt. In the second scenario, we finetuned the first three open source VMLMs on the code-mixed training set and then after combining the native language samples with the code-mixed training data.
@@ 4.1 Baseline results:
Table 2: Baselines results. Notation: originally reported scores are mentioned with ‘#\#’ mark and our reproduced results are marked as ‘†’.
@@
XLM-R $\underline{0.81}^{\#}$ ($\underline{0.83}^{\ast}$) 0.79 ($\underline{0.83}^{\ast}$) $^{\#} 0.81#0.81^{\#}$ $0.81^{\#}$ 0.80 (0.83) $0.81^{\#}$
@@
In this section, we reported our observations from the first experiment. The F1-scores obtained from all models over the considered datasets and training scenarios are reported in Table 3. The F1-scores of best-performing models for the individual training scenarios (column-wise in Table 3) were marked in bold, while the second-best results are underlined. Similarly, the best-performing scenarios giving the highest $F1$ scores for individual models (row-wise in Table 3) were kept inside the parenthesis, while the second best scores are marked with ‘#\#’ superscript. The highest score for both tasks across all training scenarios and models are marked in blue. Following are our takeaways,
  • When trained with only code-mixed samples, mBERT gave the highest F1-score of 0.78 for humour detection, followed by XLM-R and MuRIL (F1-score of 0.75). Similarly, for sarcasm detection, MuRIL reported the highest F1-score of 0.83 followed by IndicBERT and XLM-R (F1-score of 0.81).
  • On adding native samples to the code-mixed training sets, statistical classifiers did not show any performance improvement. In fact, in many cases, the performance declined sharply (a decline of 9.85% and 56.7% in F1-scores for humour and sarcasm detection, respectively).
  • Among the MLMs, mBERT showed significant improvement after native sample mixing with an F1-score of 0.84 (improvements up to 5%) in detecting sarcasm. In the case of MuRIL, we saw many instances where it resulted in statistically significant (p < 0.05) improvement (up to 5.3% and 7.2% raise in F1-score for humour and sarcasm detection respectively) after native sample mixing. Finally, we observed that IndicBERT model performed best after native sample mixing with an F1 scores of 0.79 (improvement up to 6.7%) and 0.88 (improvement up to 8.6%) in code-mixed humour and sarcasm detection respectively (both statistically significant with p< 0.05).
@@
  • We observed similar trends in precision and recall scores presented in Appendix Table 10.
@@
In this section, we presented our observations from the second experiment. The F1-scores obtained from all models for the considered datasets and training scenarios are reported in Table LABEL:tab:MTL_humor_results. In Table LABEL:tab:MTL_humor_results, each row enlists different training scenarios for considered two tasks. It has two sub-tables. The upper and lower tables report results for code-mixed humour and sarcasm detection, respectively. The first two columns in each sub-table report the combination of datasets used for training. For example, the first row under the ‘Humor’ sub-table presents the case where the training set has (i) humour samples from the code-mixed dataset and English dataset ‘ColBERT’ (written as NLD: ‘Col’), (ii) sarcasm samples from the code-mixed dataset and English dataset ‘NHD’ (tick marked next to NHD). We didn’t consider hate samples here (presented as the empty box under the ‘Hate’ column). So, we trained our models for two tasks in this case. Similarly, the fifth row represents the case where the training set has code-mixed, native English and native Hindi hate samples along with the Humour and Sarcasm samples considered in the first row. So, in this case, we trained our models for three tasks. The overall best scores were marked in blue. Following were our takeaways,
  • For code-mixed humor detection, $\text{MuRIL}_{MTL}$ reported the highest F1 score of 0.83 (up to 10.67% increment), followed by $\text{XLM-R}_{MTL}$ (0.81) and $\text{mBERT}_{MTL}$ (0.80). On the other hand, for code-mixed sarcasm detection, $\text{XLM-R}_{MTL}$ outperformed others with 0.91 F1 score (up to 12.35% increment), followed by $\text{MuRIL}_{MTL}$ (0.88) and $\text{mBERT}_{MTL}$ (0.86).
  • We achieved the highest scores in MTL strategy when native datasets with low KL divergence between classes were combined. This appears to help the pre-trained MLM-based MTL architecture focus more on contextual understanding rather than being influenced by lexical differences between the labels. Notably, SC-V2 and HCHIn, which have the lowest KL divergence, were consistently present among the best-performing configurations in both sub-tables of Table LABEL:tab:MTL_humor_results (refer [row 26, col 7] of Humor subtable and [row 32, col 5] of Sarcasm subtable).
  • For sarcasm detection, $\text{mBERT}_{MTL}$ resulted with a highest F1 score of 0.86 (improvement up to 7.5%). The improvement is statistically significant ($p<0.05$).
  • Similarly, for humor detection, $\text{XLM-R}_{MTL}$ gave an improvement of F1 score upto 0.81 (improvement up to 8%). For sarcasm detection, $\text{XLM-R}_{MTL}$ improved even more with the highest F1 score of 0.91 (improvement up to 12.35%). Both the improvements were statistically significant with $p<0.05$.
  • Finally, $\text{MuRIL}_{MTL}$ reported the highest F1 score for humor detection, i.e., 0.83 (an improvement up to 10.67%). The improvement is statistically significant ($p<0.05$). Similarly, for sarcasm detection, F1 scores ranged from 0.81 to 0.88 using $\text{MuRIL}_{MTL}$. The F1-score improved up to 6% which is statistically significant ($p<0.05$).
  • Upon analyzing the improvements deeply, we found that the MTL models performed better for samples with shorter context length (refer to Figure 5 in Appendix F) and spelling errors. Due to space constraints, we reported these insights from the ablation study in Appendix F.
  • The error analysis (refer Section 5) revealed that samples with some connection to the hate detection task improved performance in multitask setting.
  • Similar patterns appeared in the precision and recall scores (Table 11 and Table 12) as reported in the Appendix.
@@
Gemma 0.09 0.09 0.07 (0.11) $0.10^{\#}$
Aya Expanse 0.73 (0.74) $\underline{0.73}^{\#}$ 0.71 0.73
Llama-3.1 (0.75) $\textbf{0.74}^{\#}$ (0.75) (0.75) (0.75)
GPT-4 (0.74) 0.57 $0.62^{\#}$ 0.56 0.58